OpenStudy (anonymous):

How many gallons is held by a 24 inside diameter circular tank that is filled to a level of 886 inches?

OpenStudy (ranga):

No unit is specified for the diameter in the problem. Is the diameter 24 inches?

OpenStudy (anonymous):

I don't know. Thats's all I have.

OpenStudy (ranga):

The unit has got to be there. Read the problem carefully. Towards the end it says 866 inches. Similarly, next to 24 it must say 24 inches or 24 feet or some other unit..

OpenStudy (anonymous):

I'm going to send you a screenshot.

OpenStudy (ranga):

Yes, they have omitted a needed information. But we will assume it is 24 inches. Volume of a cylinder = (pi) * r^2 * h diameter = 24 inches. radius = diameter / 2 = 24/2 = 12 height h = 866. Volume = (pi) * 12^2 * 866 = ? Volume will be in cubic inches. 1 gallon = 231 cubic inches. So divide the answer you get by 231 to convert cubic inches to gallons.

OpenStudy (anonymous):

got it. Thanks